OpenStudy (binkeyboy03):

Numbers p and q are whole numbers. Which statements are always true? Choose all answers that are correct. A. p + q is a whole number B. p – q is a whole number C. p • q is a whole number D. p/q is a whole number

OpenStudy (binkeyboy03):

@YoungStudier

OpenStudy (binkeyboy03):

@KendrickLamar2014

OpenStudy (empty):

Any ideas?

OpenStudy (binkeyboy03):

no not really

OpenStudy (binkeyboy03):

@AlexandervonHumboldt2

OpenStudy (empty):

is \(\frac{1}{2}\) a whole number?

OpenStudy (binkeyboy03):

no

OpenStudy (binkeyboy03):

@AlexandervonHumboldt2

OpenStudy (binkeyboy03):

@Deadly_Roses

OpenStudy (jacobchaotic):

I'm pretty sure the answer would be A because the question says they will always be whole numbers. If you add any whole number to a whole number, you're gonna get a whole number. 2 + 2 = 4 4 + 4 = 8 6 + 6 = 12 8 + 8 = 16 10 + 10 = 20

OpenStudy (binkeyboy03):

wow thanks!

OpenStudy (binkeyboy03):

you can pick more than 1

OpenStudy (empty):

So since \(\frac{1}{2}\) isn't a whole number but 1 and 2 are whole numbers, which answer can you eliminate?